OpenStudy (anonymous):

3/4k<-6 solve the inequality

OpenStudy (anonymous):

you can flip the 3/4k to be 4k/3 and by doing that the sign will change to > so you'll have : 4k/3 > -6 = (1/3)(4k/3) > (-6)(1/3) (mulitply by 1/3 both sides to get rid of the 3) =4k > - 6 (divide by 4 to get k alone) = k> -3/2

OpenStudy (anonymous):

wait, sorry, -6/3 = -2 ; so instead the answer will be -2/4 = -1/2, sorry again :)

OpenStudy (anonymous):

wait wait, instead of multiplying by 1/3, you must multiply by 3. I apologize for this mess. so you'll have 4k > -18 = k>-9/2

myininaya (myininaya):

is the k suppose to be in the denominator?

OpenStudy (anonymous):

you can flip it to make your life easier :)

myininaya (myininaya):

is it 3k/4 or 3/(4k)?

OpenStudy (anonymous):

yor question was 3/4k >-6 right?

OpenStudy (anonymous):

if you don't flip it it's going to be: 3/k < - 24 (multiply by 4 to cancel it ) = k/3 > -24 (flip it + change signs) =k> -72 I guess.

OpenStudy (anonymous):

wait.... when you flip it, you'll have to do it to all the equation then u'll have: k/3 > - 1/24 = k>-1/8 (multiply by 3 to get rid of it) I am sooo sorry, but this is the right final answer.